find two numbers such that one of them exceeds the other by 18 and their sum is 92.
Found 2 solutions by Maths68, ssm@123:
Answer by Maths68(1474)

Let
One Number = x
Other Number = y
x=y+18 ...............(1)
x+y=92................(2)
Put the value of x from (1) to (2)
y+18+y=92
2y=92-18
2y=74
2y/2=74/2
y=37
Put the value of y in (1)
x=37+18
x=55


Check
======
Sum of numbers = 55 + 37 = 92

one Number = 55
Other Number = 37
----------------
Difference = 18
================

Answer by ssm@123(2)

Let the first number be - x
--------------------------------------
Let the second number be - x+18
--------------------------------------
the sum of the numbers is - 92
--------------------------------------
Required equation -
x+(x+18) = 92
x+x+18 = 92
2x+18 = 92
2x = 92-18
2x = 74
x = 74/2
x = 37
---------------------------------------
So,the first number - x = 37
the second number - x+18 = 37+18 = 55
=======================================
Therefore , the numbers are 37 and 55
=======================================
SOLVED BY -
SRAVYA SRI